# GmNet: Revisiting Gating Mechanisms From A Frequency View

### v1

@@ 3.2.2 Experimental Analysis
Table 1: Configurations of GmNet. We vary the embed width, the depth and the ratio to build different sizes of GmNet.
@@
4 Method
To address the limitation of low-frequency bias for CNN-based network designs, our proposed method named as GmNet integrates a simple gated linear unit into the block as illustrated in Fig. 4. GmNet offers both theoretical and practical advantages on encouraging the model to learn from a broader range of frequency regions, especially the high-frequency domain. We incorporate two depth-wise convolution layers with kernel sizes of $7\times 7$ at the beginning and end of the block respectively to facilitate the integration of low- and high-frequency information. At the core of the block, we have two $1\times 1$ convolution layers and a simple gated linear unit.
Table 2: Comparison of Efficient Models on ImageNet-1k. Latency is evaluated across various platforms, including A100 GPU and iPhone 14 mobile device. Latency benchmarking batch size is set to 1 as in real-world scenario.
@@ 4 Method
MobileV2-1.0 [23] 72.0 3.4 0.3 1.7 0.9
ShuffleV2-1.5 [17] 72.6 3.5 0.3 2.2 1.3
EfficientFormerV2-S0 [15] 73.7 3.5 0.4 2.0 0.9
StarNet-S2 [18] 74.8 3.7 0.5 1.9 0.9
@@
FasterNet-T0 [2] 71.9 3.9 0.3 2.5 0.7
EfficientFormerV2-S1 [15] 77.9 4.5 0.7 3.4 1.1
EfficientMod-xxs [19] 76.0 4.7 0.6 2.3 18.2
RepViT-M0.9 [30] 77.4 5.1 0.8 3.0 1.1
StarNet-S3 [18] 77.3 5.8 0.7 2.3 1.1
@@
EfficientMod-xs [19] 78.3 6.6 0.8 2.9 22.7
EdgeViT-XS [20] 77.5 6.8 1.2 3.0 1.6
RepViT-M1.0 [30] 78.6 6.8 1.2 3.6 1.1
StarNet-S4 [18] 78.4 7.5 1.1 3.3 1.1
Fasternet-T1 [2] 76.2 7.6 0.9 2.5 1.0
MobileOne-S2 [28] 77.4 7.8 1.3 1.8 1.0
@@
RepViT-M1.1 [30] 79.4 8.3 1.3 5.1 1.2
FastViT-S12 [27] 79.8 8.8 1.8 5.3 1.6
EfficientFormer-L1 [14] 77.2 12.3 1.3 12.1 1.4
EfficientFormerV2-S2 [15] 80.4 12.7 1.3 5.4 1.6
EfficientMod-s [19] 81.0 12.9 1.4 4.5 35.3
RepViT-M1.5 [30] 81.2 14.0 2.3 6.4 1.7
MobileOne-S4 [28] 79.4 14.8 2.9 2.9 1.8
LeViT-256 [8] 81.5 18.9 1.1 6.7 31.4
@@
We use the ReLU6 as the activation function. ReLU6 can not only emphasize the higher frequency components since it has discontinuous higher derivatives, but also it can limit the amplitudes and energy of the high-frequency components with limiting the max values. We vary the block numbers, input embedding channel numbers and channel expansion factors ‘ratio’ to build different sizes of GmNet, as detailed in Table 1.
5 Experiments
5.1 Results in Image Classification
Implementation details. We perform image classification experiments on the ImageNet-1K dataset, adopting a standard input resolution of $224\times 224$ for both training and evaluation. All model variants are trained from scratch for $300$ epochs using the AdamW optimizer, starting with an initial learning rate of $3\times 10^{-3}$ and a batch size of $2048$. The supplementary materials provide a comprehensive overview of the training setup. For performance assessment, we convert our PyTorch models into the ONNX format to measure latency on a Mobile device (iPhone 14) and a GPU (A100). Additionally, we deploy the models on the mobile device via CoreML-Tools to further evaluate latency. Importantly, our training approach does not incorporate advanced techniques such as re-parameterization or knowledge distillation. The results presented in Table 2 correspond to models trained without these enhancements.
Figure 5: Trade-off between Top-1 accuracy and latency on A100. GmNet variants achieve substantially lower latency compared to related works. A more extensive comparison is provided in the supplementary material.
Compared with the state-of-the-art. The experimental results are presented in Table 2. Without any strong training strategy, GmNet delivers impressive performance compared to many state-of-the-art lightweight models. With a comparable latency on GPU, GmNet-S1 outperforms MobileV2-1.0 by $3.5\%$. Notably, GmNet-S2 achieves $78.3\%$ with only $1.9$ms on the A100 which is a remarkable achievements for the models under 1G FLOPS. GmNet-S3 outperforms RepViT-M1.0 and StarNet-S4 by $1.9\%$ and $0.9\%$ in top-1 accuracy with $1.1$ ms and $1.4$ ms faster on the GPU latency, respectively. The improvements on the speed are over $30\%$. Additionally, with similar latency, GmNet-S3 delivers a $1.7\%$ improvement on the accuracy over MobileOne-S4. GmNet-S4 achieves 2x faster compared to RepViT-M1.5 on the GPU and it surpasses MobileOne-S4 of $2.1\%$ under the similar latencies of both GPU and Mobile. LeViT-256 [8] matches the accuracy of GmNet-S4 but runs twice as slow on a GPU and 16 times slower on an iPhone 14 The strong performance of GmNet can be largely attributed to the clear insights of gating mechanisms and simplest architectures. Fig. 5 further illustrates the latency-accuracy trade-off across different models. More comparisons and results of downstream tasks including objective detection and segmentation can be found in supplementary.
@@ 5.2 Ablation Studies
More studies on different activation functions. To further explore the effect of different activation functions, we trained various GmNet-S3 variants on ImageNet-1k. As illustrated in Fig. 4, we replaced ReLU6 with GELU, ReLU or remove the activation function. To better reflect the differences between different models, we set the radii to a larger range and only separate the decomposed image into two kinds of frequency components. For $\mathbf{z}_{i}\in(0,r)$, we recognize it as the low-frequency component while $\mathbf{z}_{i}\in(r,+\infty)$ represents the high-frequency component [32, 31]. As shown in the Table. 3, we can find that, the increases on classifying the high-frequency components are significant comparing models using and not using the activation functions. For example, comparing results of ‘Identity’ and ‘ReLU’ with the improvement of $11\%$ on the raw data, improvement on high-frequencies is over 3 times on average. ‘GELU’ and ’ReLU’ shows advances on low-/high- frequency components respectively compared to each other. This aligns with our understanding of how different types of activation functions impact frequency response. Notably, the closer performance of models with Identity and ReLU/GELU at low frequencies suggests the low-frequency bias of convolution-based networks.
@@
MobileOne-S2 [28] 77.4 35.0 11.6 6.5 36.9 2.4 53.5
EfficientMod-xs [19] 78.3 45.4 12.9 9.4 40.6 3.5 54.6
StarNet-S4 [18] 78.4 43.3 13.8 9.4 41.3 3.4 54.8
@@
Moreover, even considering the improvements on the raw data, model using the ReLU6 shows obvious increase on the high-frequency components compared to the model using GELU especially when we set $r$ to $12$, $24$, $36$. Compared to the model with ReLU, ReLU6 is more effective in preventing overfitting to high-frequency components since it has better performance on low-frequencies. Considering performances of ReLU, GELU, and ReLU6, we can observe that achieving better performance on high frequencies at the expense of lower frequencies does not necessarily lead to overall improvement, and vice versa. To get a better performance on the raw data, it is essential to enhance the model’s ability to learn various frequency signals without compromising others.
Comparison with existing methods. As addressed in Table 3, a model should achieves strong performance across different frequency components to deliver a better overall performance. However, both pure convolutional architectures and transformers exhibit a low-frequency bias, as discussed in [1, 26]. Therefore, enhancing the performance of a lightweight model depends on its ability to more effectively capture high-frequency information.
@@
To address the advantages of GmNet on overcoming the low-frequency bias, we test some existing models on different frequency components of different radii. We select three kinds of typical lightweight methods for comparison including pure conv-based model MobileOne-S2 [28], attention-based model EfficientMod-xs [19] and model also employing GLUs-like structure StarNet-S4 [18]. As shown in Table 4, accuracies of low-frequency components are close among different models considering the overall performance. However, it shows that GmNet-S3 clearly surpass the other models in high frequency components. For example, GmNet-S3 has a $6.3\%$ improvement compared to EfficientMod-xs when $r=12$ and $2.7\%$ increase when $r=24$. For StarNet, which also uses a GLU-like structure with dual-channel FC, it struggles to effectively emphasize high-frequency signals. The simplest GLUs design can deliver a better learning on various frequency components and it can achieve a better balance between the efficiency and the effectiveness.
@@
Bandwidths analysis of convolution kernels. As discussed in the [26], the convolution layer may play roles of ’smoothing’ the feature which means it has a low-frequency bias. Experiments on studying weights of the convolution layer is insightful to give more evidences of how GLUs effect the learning of different frequency components [31, 26, 1]. In this paper, we propose using the bandwidths of convolution kernels to represent their ability of responding to different frequency components. Specifically, a wider bandwidth indicates that the kernel can process a broader range of frequencies, allowing it to capture diverse frequency components simultaneously and thereby preserve rich information from the feature. As illustrated in Figure 6, the distributions of the ReLU model suggest that its convolution kernels tend to focus on a narrow range of frequency components leading to relatively lower bandwidths. It indirectly reflects an overemphasis on high-frequency components. Although the model using GELU exhibits a better distribution in the top convolutional layers, it still has a low-frequency bias, leading to a distribution shift in the bottom convolutional layers. Compared to other activations, the enhanced bandwidth distribution of the model using ReLU6 demonstrates better generalization for this task. The properties of the convolution kernels align with the results in Table 3.
@@ 6 Conclusion
In this paper, we systematically explore the effect of gating mechanisms on the training dynamics of neural networks from a frequency perspective. We provide both theoretical and experimental analysis to reveal how the element-wise product and activation interplay to manage the balance between different frequencies learning. The element-wise product can broaden the spectrum of the signals while introduce high-frequency noise. The activation function can effectively control the frequency output based on its own characters. Benefited by the comprehensive study on GLUs, we introduce a lightweight model named GmNet to efficiently utilize the information of various frequency. It shows impressive performance on both effectiveness and efficiency compared with recent lightweight models.

### v2

@@ 5.1 Results in Image Classification
Table 1: Configurations of GmNet. We vary the embedding width, depth, and gating ratio to construct different model sizes of GmNet.
@@
Figure 6: Trade-off between Top-1 accuracy and latency on A100. GmNet variants achieve substantially lower latency compared to related works, while maintaining competitive or superior Top-1 accuracy. This demonstrates the efficiency of our design in balancing performance and computational cost. A more extensive comparison is provided in the supplementary material.
Table 2: Comparison of Efficient Models on ImageNet-1k. Latency is evaluated across various platforms (A100 GPU and iPhone 14).
@@
FasterNet-T0 chen2023run 71.9 3.9 0.3 2.5 0.7
MobileV2-1.0 sandler2018mobilenetv2 72.0 3.4 0.3 1.7 0.9
ShuffleV2-1.5 ma2018shufflenet 72.6 3.5 0.3 2.2 1.3
EfficientFormerV2-S0 li2023rethinking 73.7 3.5 0.4 2.0 0.9
MobileNetv4-Conv-S qin2024mobilenetv4 73.8 3.8 0.2 2.2 0.9
StarNet-S2 ma2024rewrite 74.8 3.7 0.5 1.9 0.9
LSNet-T wang2025lsnet 74.9 11.4 0.3 2.9 1.8
@@
EfficientMod-xxs ma2024efficient 76.0 4.7 0.6 2.3 18.2
Fasternet-T1 chen2023run 76.2 7.6 0.9 2.5 1.0
EfficientFormer-L1 li2022efficientformer 77.2 12.3 1.3 12.1 1.4
StarNet-S3 ma2024rewrite 77.3 5.8 0.7 2.3 1.1
MobileOne-S2 vasu2023mobileone 77.4 7.8 1.3 1.9 1.0
RepViT-M0.9 wang2024repvit 77.4 5.1 0.8 3.0 1.1
EfficientFormerV2-S1 li2023rethinking 77.9 4.5 0.7 3.4 1.1
@@
EfficientMod-xs ma2024efficient 78.3 6.6 0.8 2.9 22.7
StarNet-S4 ma2024rewrite 78.4 7.5 1.1 3.3 1.1
SwiftFormer-S shaker2023swiftformer 78.5 6.1 1.0 3.8 1.1
RepViT-M1.0 wang2024repvit 78.6 6.8 1.2 3.6 1.1
UniRepLKNet-F ding2024unireplknet 78.6 6.2 0.9 3.1 3.5
@@
RepViT-M1.1 wang2024repvit 79.4 8.3 1.3 5.1 1.2
MobileOne-S4 vasu2023mobileone 79.4 14.8 2.9 2.9 1.8
FastViT-S12 vasu2023fastvit 79.8 8.8 1.8 5.3 1.6
MobileNetv4-Conv-M qin2024mobilenetv4 79.9 9.2 1.0 9.2 1.4
LSNet-B wang2025lsnet 80.3 23.2 1.3 6.2 3.6
EfficientFormerV2-S2 li2023rethinking 80.4 12.7 1.3 5.4 1.6
EfficientMod-s ma2024efficient 81.0 12.9 1.4 4.5 35.3
RepViT-M1.5 wang2024repvit 81.2 14.0 2.3 6.4 1.7
LeViT-256 graham2021levit 81.5 18.9 1.1 6.7 31.4
@@ 5.2 Ablation Studies
More studies on different activation functions. To further explore the effect of different activation functions, we trained various GmNet-S3 variants on ImageNet-1k. As illustrated in Fig. 5, we replaced ReLU6 with GELU, ReLU or remove the activation function. To better reflect the differences between different models, we set the radii to a larger range/
@@
As shown in the Table. 3, we can find that, the increases on classifying the high-frequency components are significant comparing models using and not using the activation functions. For example, comparing results of ‘Identity’ and ‘ReLU’ with the improvement of $11\%$ on the raw data, improvement on high-frequencies is over 3 times on average. ‘GELU’ and ’ReLU’ shows advances on low-/high- frequency components respectively compared to each other. This aligns with our understanding of how different types of activation functions impact frequency response. Notably, the closer performance of models with Identity and ReLU/GELU at low frequencies suggests the low-frequency bias of convolution-based networks.
Moreover, even considering the improvements on the raw data, model using the ReLU6 shows obvious increase on the high-frequency components compared to the model using GELU especially when we set $r$ to $12$, $24$, $36$. Compared to the model with ReLU, ReLU6 is more effective in preventing overfitting to high-frequency components since it has better performance on low-frequencies. Considering performances of ReLU, GELU, and ReLU6, we can observe that achieving better performance on high frequencies at the expense of lower frequencies does not necessarily lead to overall improvement, and vice versa. To get a better performance on the raw data, it is essential to enhance the model’s ability to learn various frequency signals.
Comparison with existing methods. As addressed in Table 3, a model should achieves strong performance across different frequency components to deliver a better overall performance.
@@
MobileOne-S2 vasu2023mobileone 77.4 35.0 11.6 6.5 36.9 2.4 53.5
EfficientMod-xs ma2024efficient 78.3 45.4 12.9 9.4 40.6 3.5 54.6
StarNet-S4 ma2024rewrite 78.4 43.3 13.8 9.4 41.3 3.4 54.8
@@
However, both pure convolutional architectures and transformers exhibit a low-frequency bias, as discussed in bai2022improving ; tang2022defects . Therefore, enhancing the performance of a lightweight model depends on its ability to more effectively capture high-frequency information. To address the advantages of GmNet on overcoming the low-frequency bias, we test some existing models on different frequency components of different radii. We select three kinds of typical lightweight methods for comparison including pure conv-based model MobileOne-S2 vasu2023mobileone , attention-based model EfficientMod-xs ma2024efficient and model also employing GLUs-like structure StarNet-S4 ma2024rewrite . As shown in Table 4, accuracies of low-frequency components are close among different models considering the overall performance. However, it shows that GmNet-S3 clearly surpass the other models in high frequency components. For example, GmNet-S3 has a $6.3\%$ improvement compared to EfficientMod-xs when $r=12$ and $2.7\%$ increase when $r=24$. For StarNet, which also uses a GLU-like structure with dual-channel FC, it struggles to effectively emphasize high-frequency signals. The simplest GLUs design can achieve a better balance between the efficiency and the effectiveness.
@@
Bandwidths analysis of convolution kernels. As discussed in the tang2022defects , the convolution layer may play roles of ’smoothing’ the feature which means it has a low-frequency bias. Experiments on studying weights of the convolution layer is insightful to give more evidences of how GLUs effect the learning of different frequency components wang2020high ; tang2022defects ; bai2022improving . In this paper, we propose using the bandwidths of convolution kernels to represent their ability of responding to different frequency components. Specifically, a wider bandwidth indicates that the kernel can process a broader range of frequencies, allowing it to capture diverse frequency components simultaneously and thereby preserve rich information from the feature. As illustrated in Figure 7, the distributions of the ReLU model suggest that its convolution kernels tend to focus on a narrow range of frequency components leading to relatively lower bandwidths. It indirectly reflects an overemphasis on high-frequency components. Although the model using GELU exhibits a better distribution in the top convolutional layers, it still has a low-frequency bias, leading to a distribution shift in the bottom convolutional layers. Compared to other activations, the enhanced bandwidth distribution of the model using ReLU6 demonstrates better generalization for this task. The properties of the convolution kernels align with the results in Table 3.
@@
6 Limitations
Discussions. Different datasets and training strategies may lead to different dynamics which has been addressed in wang2020high ; tang2022defects . For the small dataset which is used in Sec. 3, the convergence is faster which means NNs can easily capture some certain frequency components and then learning from other frequency components becomes very slow. Moreover, from the performance on very high-frequency components (for $r>36$), it is evident that neural networks face challenges in directly classifying these components, which are nearly imperceptible to the human eye. However, NNs can be sensitive to the perturbation of high-frequency noise yin2019fourier . This is also an interesting topic of revealing the robustness of NNs, but it is beyond the scope of this article.
Limitations. This work investigates that enhancing high-frequency learning is crucial for improving overall model performance for efficient models. However, this does not imply that GLUs are a universally optimal solution for all architectures. Larger models, with higher embedding dimensions and deeper structures, may inherently mitigate low-frequency bias, reducing the relative impact of such gating mechanisms. In this paper, we focus specifically on efficient networks. Additionally, in architectures with more complex structures, simply replacing the activation function with a GLU may not significantly enhance the model’s ability to capture high-frequency information.
@@ 7 Conclusion
In this paper, we systematically explore the effect of gating mechanisms on the training dynamics of neural networks from a frequency perspective. We provide both theoretical and experimental analysis to reveal how the element-wise product and activation interplay to manage the balance between different frequencies learning. The element-wise product can broaden the spectrum of the signals while introduce high-frequency noise. The activation function can effectively control the frequency output based on its own characters. Benefited by the comprehensive study on GLUs, we introduce a lightweight model named GmNet to efficiently utilize the information of various frequency. It shows impressive performance on both effectiveness and efficiency.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state our contributions and scope in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in Sec. 6
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We have provided the full set of assumptions and a complete (and correct) proof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have included all details of our implementation both in main text and appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: The data used in this paper are all public and we will release our code upon acceptance
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have included details of our implementation in Sec. 5
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We cover the details in Sec. 5 about evaluation setups.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided details of our compute resources in the Sec. 5.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we follow all the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: They are in Sec. 6
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work has no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the original papers of the datasets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We do not use LLM for help.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.